[INSTRM-571] Restore H4RG readouts with new cabling. Created: 15/Nov/18 Updated: 17/Nov/18 Resolved: 17/Nov/18 |
|
| Status: | Done |
| Project: | Instrument control development |
| Component/s: | None |
| Affects Version/s: | None |
| Fix Version/s: | None |
| Type: | Task | Priority: | Blocker |
| Reporter: | cloomis | Assignee: | cloomis |
| Resolution: | Done | Votes: | 0 |
| Labels: | SM1 | ||
| Remaining Estimate: | Not Specified | ||
| Time Spent: | Not Specified | ||
| Original Estimate: | Not Specified | ||
| Description |
|
There are two built up H4 signal chains at JHU: one in the n1 cryostat in the NIR cleanroom, and one in n9 in the Optics lab. Both have new flight cabling: the SAM in is a vacuum feedthrough box, the braided cable harness to the ASIC at the FPA, and the proper ASIC-FPA ribbon. All parts have been tested separately and together. But the n9 reads currently look like there is no detector. More specifically, they look like the ASIC is always configured with grounded inputs (bypassing the inputs from the detector). This is a warning ticket, without specific work or times. |
| Comments |
| Comment by cloomis [ 17/Nov/18 ] |
|
Damaged connector on the ASIC, for the detector cable. ASIC and cable replaced. |